# Language all the way down: Grammatical structures in mathematics

### v2

@@ Introduction
In Equation (1) we also see an example of 2 becoming un-nominalised, returning to quantifier form. The construction of $2x$ as a noun group uses 2 as a quantifier describing the noun $x$. Herb and Ken Gross are credited with an adjective-noun (technically quantifier-noun) description of these algebraic objects many years ago (the adjective-noun algebra description has inspired many people as evidenced by the material on the now-defunct website adjectivenounmath.com [5]) but I disagree with their claim that arithmetic alone without variables (such as) $2+1=3$ has a hidden common noun that the numbers are attached to as quantifiers. Instead I think that 1, 2, and 3 are present as nouns from the nominalisation process, and it is plausible that one of the difficulties of introducing the quantifier-noun algebraic construction is undoing the nominalisation process. We go from $2\times x$ where both 2 and $x$ are present as nouns to $2x$ – the quantifier-noun group construction that has been transformed grammatically and mathematically by the now hidden multiplication process. We still have 1 appearing as a noun though, so the student has to hold both grammatical forms in play simultaneously.
In the first example we have introduced some linguistic terminology in relation to a simple function definition. From here we are going to move into a far more complicated function that is more directly relevant to teaching statistics – the normal distribution probability density function.
@@
As in the first example, our naming of the probability density function $f(x)$ is instantiation as a noun group. I read $=$ as ‘is equal to’, which combines the verb ‘is’ and the prepositional phrase ‘equal to’, where the prepositional phrase is used to express a relationship, in this case mathematical equality between the two sides.
@@
Next we have the exponential function and its argument, which is combined with the first part through implicit multiplication. As with the other functions, $\exp$ can be understood as a modifier on the noun group argument. The argument in this case is quite complicated. We start with -- as a negative modifier on the fraction. The numerator has a power modifying the phrase $x-\mu$, which is a second use of -- but this time as a verb combining the nouns $x$ and $\mu$. The dual use of -- as an arithmetic operation and a valence modifier, one of which takes two arguments and the other one, can be unhelpful. The fraction line represents division as a verb, and the denominator is another noun phrase constructed through multiplication of noun $2$ and noun phrase $\sigma^{2}$.
@@
The expression in Equation (3) has a near-identical grammatical construction to Equation (1): on the left side a noun $Y$, on the right hand a noun group $X\beta$ and noun $\epsilon$ combined by the verb $+$. However, we have changed the meaning by sneakily swapping to matrix rather than number arithmetic. In the matrix form, it is less appropriate to read $\beta$ as a modifier on $X$, but rather as a noun in its own right. The implicit change in operations to matrix arithmetic can trip students up because we use the same arithmetic symbols and grammatical construction – we introduce lexical ambiguity through multiply defined symbols. We can still read $Y$, $X$, and $\epsilon$ as nouns, and $+$ as a verb, but there is a considerably greater amount of information tied up in the symbols compared to Equation (1) because the data contained in $X$ is no longer a single number, and the operations of matrix arithmetic are combinatorial number arithmetic rather than a single instance thereof. The neutron-star like density of information of Equation (3) is precisely why we don’t require any more symbols to express the relationship, but comes at the cost of making the matrix operation structure implicit.
@@
Equations (3) and (4), combined with the initial naming of the variables and statement that they have a linear relationship, gives us three different representations of the same underlying structure. There are differing levels of abstraction depending on how explicit the number arithmetic is made. The term ‘linear model’ is the most abstract, and is the nominalisation of the overall relationship.
@@
The core argument of this article is that making linguistic structure clear to students may help them leverage their existing language skills to understand mathematical objects and representations. I have explicitly taught these structures in my classes and found it useful, but there is no literature testing the results as the linguistic structures I am discussing are not part of existing language theory. However, I can offer some suggestions for how to introduce the grammatical structure I discuss.

### v3

# Grammatical structures in mathematics: a personal view

@@ Introduction
In Equation (1) we also see an example of 2 becoming un-nominalised, returning to quantifier form. The construction of $2x$ as a noun group uses 2 as a quantifier describing the noun $x$. For languages other than English that use a morphological structure for quantifiers, this may be more directly compared to a quantifier prefix 2 and a noun root $x$ to give the compound $2x$. Herb and Ken Gross are credited with an adjective-noun (technically quantifier-noun) description of these algebraic objects many years ago (the adjective-noun algebra description has inspired many people as evidenced by the material on the now-defunct website adjectivenounmath.com [5]) but I disagree with their claim that arithmetic alone without variables (such as)
$2+1=3$ (2)
has a hidden common noun that the numbers are attached to as quantifiers. Instead I think that in Equation (2), 1, 2, and 3 are present as nouns, and it is plausible that one of the difficulties of introducing the quantifier-noun algebraic construction is undoing the nominalisation process. For comparison, in Equation (1) we go from $2\times x$ where both 2 and $x$ are present as nouns to $2x$ – the quantifier-noun group construction that has been transformed grammatically and mathematically by the now hidden multiplication process. We still have 1 appearing as a noun though, so the student has to hold both grammatical forms in play simultaneously.
In the first example we have introduced some linguistic terminology in relation to a simple function definition. From here we are going to move into a far more complicated function that is more directly relevant to teaching statistics – the normal distribution cumulative distribution function.
@@
As in the first example, our naming of the cumulative distribution function $f(x)$ is instantiation as a noun group. I read $=$ as ‘is equal to’, which combines the verb ‘is’ and the prepositional phrase ‘equal to’, where the prepositional phrase is used to express a relationship, in this case mathematical equality between the two sides.
@@
Next we have the exponential function and its argument, which is combined with the first part through implicit multiplication. As with the other functions, $\exp$ can be understood as a modifier on the noun group argument. The argument in this case is quite complicated. We start with -- as a negative modifier on the fraction. The numerator has a power modifying the phrase $x-\mu$, which is a second use of -- but this time as a verb combining the nouns $x$ and $\mu$. The dual use of -- as an arithmetic operation and a valence modifier indicating a negative number, the first of which takes two arguments and the second one, can be unhelpful. The fraction line represents division as a verb, and the denominator is another noun phrase constructed through multiplication of noun $2$ and noun phrase $\sigma^{2}$.
@@
The expression in Equation (4) has a near-identical grammatical construction to Equation (1): on the left side a noun $Y$, on the right hand a noun group $X\beta$ and noun $\epsilon$ combined by the verb $+$. However, we have replaced the operations with matrix rather than number arithmetic. In the matrix form, it is less appropriate to read $\beta$ as a modifier on $X$, but rather as a noun in its own right. The implicit change in operations to matrix arithmetic can trip students up because we use the same arithmetic symbols and grammatical construction – we introduce lexical ambiguity through multiply defined symbols. We can still read $Y$, $X$, and $\epsilon$ as nouns, and $+$ as a verb, but there is a considerably greater amount of information tied up in the symbols compared to Equation (1) because the data contained in $X$ is no longer a single number, and the operations of matrix arithmetic are combinatorial number arithmetic rather than a single instance thereof. The high density of information of Equation (4) is precisely why we don’t require any more symbols to express the relationship, but comes at the cost of making the matrix operation structure implicit.
@@
Equations (4) and (5), combined with the initial naming of the variables and statement that they have a linear relationship, gives us three different representations of the same underlying structure. There are differing levels of abstraction depending on how explicit the number arithmetic is made. The term ‘linear model’ is the most abstract, and is the nominalisation of the overall relationship. I think part of mathematical fluency is the ability to move between different representations of the same object depending on which highlights relevant structure at the time, a simple example of which would be going back and forth with summation notation such as
$1+2+3+4+\ldots+n=\sum_{i=1}^{n}i.$
@@
The core argument of this article is that making linguistic structure clear to students may help them leverage their existing language skills to understand mathematical objects and representations. Existing skills may include formal grammar training in schools, which makes it easier to introduce terminology such as nouns and verbs, but in the absence of language structure education it may still be possible to use direct comparison to phrases with similar construction. For example $2x$ is comparable to ‘two apples’, which I have observed in mathematics classrooms at the high school level but without the explicit intent of leveraging grammatical structure.
As statisticians, we use mathematical objects as models, and rely on mathematical language to describe them, which can be a barrier to statistical users and students who do not have the opportunity to develop good mathematical fluency. Understanding language structure in mathematical expressions and making them explicit may help overcome such a barrier. How mathematical language is used in the specific context of statistics may be different to, for example, theoretical mathematics.
I have explicitly taught these grammatical structures in my classes and found it useful, but there is no literature testing the results as the linguistic structures I am discussing are not part of existing language theory. However, I can offer some suggestions for how to introduce the grammatical structure I discuss.
@@
The ideas presented in this article are based on English mathematical reading specifically, but many of the same structures are likely transferable, such as nouns and quantifiers. In a mixed language classroom, inviting students to map out grammatical structures across languages may reinforce learning, but would take time and effort that is rarely available to teachers.
@@
Due to the absence of existing work, there is a fertile field of research available in the linguistics of mathematics, open to linguists, mathematicians, and statisticians alike. The initial step would be to develop a grammar of mathematical expressions. Classroom experiments could test whether that grammatical structure is useful for teaching. For statisticians, an explicit grammar is a tool for understanding how we use mathematical objects as models, and talk about the relationship between those objects, data, and the underlying thing we are trying to model. We can leverage such a grammar to help us communicate these relationships to non-experts and students alike.